# EW D-optimal Designs for Experiments with Mixed Factors

### v1

@@ 4 Applications to Real Experiments
In this section, we use the paper feeder experiment (see Section 1) and a minimizing surface defects experiment, both under MLMs, to illustrate the advantages gained by adopting our algorithms and the EW D-optimal designs. For examples under GLMs, please see Section S7 of the Supplementary Material.
@@ 4.1 Paper feeder experiment
In this section, we consider the motivating example mentioned in Section 1, the paper feeder experiment. There are eight discrete control factors and one continuous factor, the stack force, under our consideration (see Section S5 for more details). As mentioned in Section 1, instead of using two separate GLMs to model the two types of failures, misfeed and multifeed, we consider multinomial logistic models (MLM) to model all three possible outcomes, namely misfeed, normal, and multifeed, at the same time. Using Akaike information criterion (AIC, Akaike (1973)) and Bayesian information criterion (BIC, Hastie et al. (2009)), we adopt a cumulative logit model with npo as the most appropriate model for this experiment (see Table S.1 in Section S5 of the Supplementary Material).
@@
For this experiment, we compare ten different designs (see Tables 1 and 2): (1) “Original allocation”: the original design that collected 1,785 observations roughly uniformly at 183 distinct experimental settings; (2) “EW Bu-appro”: an approximate design obtained by applying Bu et al. (2020)’s EW lift-one algorithm on the original 183 distinct settings; (3) “EW Bu-exact”: an exact design obtained by applying Bu et al. (2020)’s EW exchange algorithm on the 183 distinct settings with $n=1,785$; (4) “EW Bu-grid2.5” & (5) “EW Bu-grid2.5 exact”: approximate and exact designs by applying Bu et al. (2020)’s algorithms after discretizing the stack force factor in $[0,160]$ into grid points with equal space 2.5; (6) “EW ForLion”: the proposed EW D-optimal approximate design by applying our EW ForLion algorithm to the continuous factor in $[0,160]$; (7$\sim$(10) exact designs obtained by applying Algorithm 2 with different grid levels or pace lengths ($L=0.1,0.5,1,2.5$) and $n=1,785$. For illustration purpose, all designs follow the same 18 runs modified from $OA(18,2^{1}\times 2^{7})$ for the eight discrete factors, while their design points can be different in terms of the continuous factor. Since the feasible space of a cumulative logit model is not rectangular (see Example 5.2 in Bu et al. (2020) or Example 8 in Huang et al. (2025)), we bootstrap the original dataset to collect $B=100$ (due to computational intensity) samples and the corresponding estimated model parameters, denoted by $\hat{\boldsymbol{\theta}}_{1},\hat{\boldsymbol{\theta}}_{2},\ldots,\hat{\boldsymbol{\theta}}_{100}$ All the EW designs are under the sample-based EW D-optimality with respect to the 100 parameter vectors. The constructed designs are presented in Tables LABEL:tab:D-Optimal_designs_for_the_paper_feeder_experiment and LABEL:tab:EW_Forlion_optimal_designs of the Supplementary Material.
@@
In Table 1, we list the number of design points $m$, the computational time in seconds for obtaining the designs, the objective function values, and the relative efficiencies compared with the recommended EW ForLion design, that is, $(|\hat{E}\{{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\Theta})\}|/|\hat{E}\{{\mathbf{F}}(\boldsymbol{\xi}_{\rm ForLion},$ $\boldsymbol{\Theta})\}|)^{1/p}$ with $p=32$. Compared with our EW ForLion design, the original design on 183 settings only attains $63.03\%$ relative efficiency. Having applied Bu et al. (2020)’s algorithms on the same 183 settings, the relative efficiencies of EW Bu-appro and EW Bu-exact, $88.16\%$ and $88.15\%$, are not satisfactory either. When we apply Bu et al. (2020)’s algorithms on grid points with length $2.5$, it takes much longer time to find EW Bu-grid2.5 and EW Bu-grid2.5 exact designs, which also need more settings (55 and 44). The EW ForLion design and its exact designs at different grid levels require the minimum number of experimental settings, shorter computational time, and higher relative efficiencies.
@@
Table 2: Relative efficiencies w.r.t. locally D-optimal designs for paper feeder experiment
@@ 4.2 Minimizing surface defects
Wu (2008) presented a study that investigated optimal parameter settings in a polysilicon deposition process for circuit manufacturing. This experiment involves responses of five categories, five continuous control factors, and one discrete control factor (see Table S1 in the Supplementary Material of Huang et al. (2024) or Table 6 in Lukemire et al. (2022)). Lukemire et al. (2022) employed a PSO algorithm and derived a locally D-optimal approximate design with 14 support points under a cumulative logit model with po, while Huang et al. (2024) applied their ForLion algorithm and obtained a locally D-optimal approximate design with 17 design points.
In practice, an experimenter may not know the precise nominal values for some or all parameters but often has varying degrees of insight into their actual values. To construct a robust design against parameter misspecifications, we sample $B=1000$ parameter vectors from the prior distributions listed in Table S1 of the Supplementary Material of Huang et al. (2024) for the cumulative logit model with po as follows:
@@
with $i=1,\ldots,m$, $j=1,2,3,4$, and $J=5$, where $\pi_{ij}\in(0,1)$ stands for the probability that the response associated with ${\mathbf{x}}_{i}=(x_{i1},\ldots,x_{i6})^{T}$ falls into the $j$th category.
@@
To compare our EW ForLion D-optimal designs with Bu et al. (2020)’s EW D-optimal designs constructed on grid points, we discretize each of the five continuous control factors into 2 or 4 evenly distributed grid points, which leads to $642\times 2^{5}=$ 64 or $0482\times 4^{5}=2,$ 048 possible design points, respectively. Discretizations with higher number of grid points are skipped due to computational intensity. Then Bu et al. (2020)’s lift-one algorithm is applied to the discretized design region and the corresponding EW D-optimal approximate designs are constructed (EW Bu grid2 and EW Bu grid4 in Figure 1).
@@
Design Cleaning Design Cleaning
point Method Temp. Pressure Nit.Flow Sil.Flow Time $w_{i}$ point Method Temp. Pressure Nit.Flow Sil.Flow Time $n_{i}$
@@
Figure 1: Relative efficiency comparison among designs for 10000 sampled parameter sets in the minimizing surface defects experiment
@@ 5 Conclusion
First-order Derivative of Sensitivity Function
@@
Proofs of Main Theorems
@@
Model Selection and Design Comparison for Paper Feeder Experiment
@@
In Bu et al. (2020), it was mentioned that (S.1) has analytic solutions for $J\leq 5$, while no formula was provided accordingly. In this section, we provide explicit solutions to facilitate the programming for multinomial logistic models. Note that we only need the real-valued solutions locating in $(0,1)$ for our purposes.
@@
Lemma S1.1.
When $J=3$, (S.1) can be rewritten as $A_{2}z^{2}+A_{1}z+A_{0}=0$ with $A_{2}=p(b_{0}-b_{1}+b_{2})$, $A_{1}=b_{1}(1+p)-2(b_{0}p+b_{2})$, and $A_{0}=b_{0}p-b_{1}$
@@
    If $A_{2}=A_{1}=A_{0}=0$, then (S.1) has infinitely many solutions in $(0,1)$ and $f_{i}(z)\equiv f_{i}(0)\geq 0$. In this case, we may choose $z=0$ to maximize $f_{i}(z)$.
@@
    If $A_{2}=A_{1}=0$ but $A_{0}\neq 0$, then we must have $A_{0}>0$. In this case, (S.1) has no solution, while $z=0$ uniquely maximizes $f_{i}(z)$ with $z\in[0,1]$.
@@
    If $A_{2}=0$ but $A_{1}\neq 0$, then (S.1) has a unique solution $-A_{0}/A_{1}\in\mathbb{R}$.
@@
    If $A_{2}\neq 0$ and $A_{1}^{2}-4A_{2}A_{0}>0$, then (S.1) has two real solutions
@@
    If $A_{2}\neq 0$ and $A_{1}^{2}-4A_{2}A_{0}=0$, then (S.1) has only one real solution $-A_{1}/(2A_{2})$.
@@
    If $A_{2}\neq 0$ and $A_{1}^{2}-4A_{2}A_{0}<0$, then (S.1) has no real solution.
Proof of Lemma S1.1: When $J=3$, (S.1) can be rewritten as $A_{2}z^{2}+A_{1}z+A_{0}=0$ with $A_{2}=p(b_{0}-b_{1}+b_{2})$, $A_{1}=b_{1}(1+p)-2(b_{0}p+b_{2})$, $A_{0}=b_{0}p-b_{1}$ , and
@@
Lemma S1.2.
When $J=4$, equation (S.1) is equivalent to $A_{3}z^{3}+A_{2}z^{2}+A_{1}z+A_{0}=0$ with $A_{0}=b_{0}p-b_{1}$ , $A_{1}=-3b_{0}p+b_{1}(2+p)-2b_{2}$ , $A_{2}=3b_{0}p-b_{1}(1+2p)+b_{2}(2+p)-3b_{3}$ , and $A_{3}=p(-b_{0}+b_{1}-b_{2}+b_{3})$. The cases with $A_{3}=0$ has been listed in Lemma S1.1. If $A_{3}\neq 0$, (S.1) is equivalent to $z^{3}+a_{2}z^{2}+a_{1}z+a_{0}=0$ with $a_{i}=A_{i}/A_{3}$, $i=0,1,2$. We let
@@
    If $q^{3}+r^{2}>0$, then (S.1) has only one real solution $s_{1}+s_{2}-a_{2}/3$
@@
    If $q^{3}+r^{2}=0$, then (S.1) has two real solutions $z_{1}=2r^{1/3}-a_{2}/3$ and $z_{2}=-r^{1/3}-a_{2}/3$
@@
    If $q^{3}+r^{2}<0$, then (S.1) has three real solutions
@@
Lemma S1.3.
When $J=5$, equation (S.1) is equivalent to $A_{4}z^{4}+A_{3}z^{3}+A_{2}z^{2}+A_{1}z+A_{0}=0$ with $A_{0}=b_{0}p-b_{1}$ , $A_{1}=-4b_{0}p+b_{1}(3+p)-2b_{2}$ , $A_{2}=6b_{0}p-3b_{1}(1+p)+b_{2}(4+p)-3b_{3}$ , and $A_{3}=-4b_{0}p+b_{1}(1+3p)-2b_{2}(1+p)+b_{3}(3+p)-4b_{4}$ , and $A_{4}=p(b_{0}-b_{1}+b_{2}-b_{3}+b_{4})$. The cases with $A_{4}=0$ has been listed in Lemmas S1.2 & S1.1. If $A_{4}\neq 0$, (S.1) is equivalent to $z^{4}+a_{3}z^{3}+a_{2}z^{2}+a_{1}z+a_{0}=0$ with $a_{i}=A_{i}/A_{4}$, $i=0,1,2,3$. Then there are four solutions to (S.1) calculated as complex numbers:
@@
Following Fedorov and Leonov (2014) and Huang et al. (2024), in this section we list the assumptions needed for this paper. Note that our assumptions have been adjusted for mixed factors. Recall that a design point or experimental setting ${\mathbf{x}}=(x_{1},\ldots,x_{d})^{T}\in{\cal X}\subseteq\mathbb{R}^{d}$. Among the $d\geq 1$ factors, the first $k$ factors are continuous, and the last $d-k$ factors are discrete. Note that $0\leq k\leq d$ and $k=0$ implies no continuous factor. When $k\geq 1$, we denote ${\mathbf{x}}_{1}=(x_{1},\ldots,x_{k})^{T}$ as the vector of continuous factors.
  • (A1)
    The design region ${\cal X}\subset\mathbb{R}^{d}$ is compact.
    (In this paper, ${\cal X}=\prod_{j=1}^{d}I_{j}$ , where $I_{j}=[a_{j},b_{j}]$ with $-\infty<a_{j}<b_{j}<\infty$ for continuous factors, and $I_{j}=\{c_{j1},\ldots,c_{jm_{j}}\}$ with $-\infty<a_{j}=c_{j1}<\cdots<c_{jm_{j}}=b_{j}<\infty$ and $m_{j}\geq 2$ for discrete factors.)
  • (A2)
    For each $\boldsymbol{\theta}\in\boldsymbol{\Theta}$, if $k\geq 1$, the Fisher information ${\mathbf{F}}({\mathbf{x}},\boldsymbol{\theta})$ is element-wise continuous with respect to all continuous factors of ${\mathbf{x}}\in{\cal X}$.
  • (A3)
    For each $\boldsymbol{\theta}\in\boldsymbol{\Theta}$, if $k\geq 1$, the Fisher information $\mathbf{F}(\mathbf{x},\boldsymbol{\theta})=(F_{st}({\mathbf{x}},\boldsymbol{\theta}))_{s,t=1,\ldots,p}$ is element-wise continuous with respect to all continuous factors of ${\mathbf{x}}\in{\cal X}$, and there exists a nonnegative and integrable function $K(\boldsymbol{\theta})$ on $\boldsymbol{\Theta}$, such that, $\int_{\boldsymbol{\Theta}}K(\boldsymbol{\theta})Q(d\boldsymbol{\theta})<\infty$, and $|F_{st}(\mathbf{x},\boldsymbol{\theta})|\leq K(\boldsymbol{\theta})$, for all $s,t\in\{1,\ldots,p\}$, $\mathbf{x}\in\mathcal{X}$, and $\boldsymbol{\theta}\in\boldsymbol{\Theta}$.
Note that (i) if $k=0$, that is, all factors are discrete, then both Assumption (A2) and the first half (i.e., the continuous statement) of Assumption (A3) are automatically satisfied; (ii) when $k\geq 1$, that is, there is at least one continuous factor, (A3) always implies (A2).
Lemma S2.1.
@@
The proof of Lemma S2.1 is straightforward. It is for sample-based EW designs. For integral-based EW designs, we need the following lemma:
Lemma S2.2.
@@
For any $s,t\in\{1,\ldots,p\}$, since $F_{st}(\mathbf{x},\boldsymbol{\theta})$ is continuous with respect to $\mathbf{x}_{1}=(x_{1},\ldots,x_{k})^{T}$, we must have $\lim_{n\rightarrow\infty}F_{st}({\mathbf{x}}_{n},\boldsymbol{\theta})=F_{st}({\mathbf{x}}_{0},\boldsymbol{\theta})$ for each $\boldsymbol{\theta}\in\boldsymbol{\Theta}$. Since (A3) is satisfied, according to the Dominated Convergence Theorem (see, for example, Theorem 5.3.3 in Resnick (2003)), we must have
$\lim_{n\rightarrow\infty}E\{F_{st}({\mathbf{x}}_{n},\boldsymbol{\Theta})\}=\lim_{n\rightarrow\infty}\int_{\boldsymbol{\Theta}}F_{st}({\mathbf{x}}_{n},\boldsymbol{\theta})Q(d\boldsymbol{\theta})=\int_{\boldsymbol{\Theta}}F_{st}({\mathbf{x}}_{0},\boldsymbol{\theta})Q(d\boldsymbol{\theta})=E\{F_{st}({\mathbf{x}}_{0},\boldsymbol{\Theta})\}$
for each ${\mathbf{x}}_{0}\in{\cal X}$, as long as $\lim_{n\rightarrow\infty}{\mathbf{x}}_{n}={\mathbf{x}}_{0}$ That is, $E\{{\mathbf{F}}({\mathbf{x}},\boldsymbol{\Theta})\}$ is element-wise continuous with respect to ${\mathbf{x}}_{1}$ ∎
Following Fedorov and Leonov (2014), to characterize the sample-based and integral-based D-optimal EW designs, we extend the collection $\boldsymbol{\Xi}$ of designs (each design consists of only a finite number of design points) to $\boldsymbol{\Xi}({\cal X})$, which consists of all probability measures on ${\cal X}$. In other words, a design $\boldsymbol{\xi}\in\boldsymbol{\Xi}({\cal X})$ is also a probability measure on ${\cal X}$, and for each $\boldsymbol{\theta}\in\boldsymbol{\Theta}$,
Then for sample-based EW criterion, we denote
${\mathbf{F}}_{\rm SEW}(\boldsymbol{\xi})=\hat{E}\{{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\Theta})\}=\frac{1}{B}\sum_{j=1}^{B}{\mathbf{F}}(\boldsymbol{\xi},\hat{\boldsymbol{\theta}}_{j})=\int_{\cal X}\hat{E}\{{\mathbf{F}}({\mathbf{x}},\boldsymbol{\Theta})\}\boldsymbol{\xi}(d{\mathbf{x}})\ .$ (S.2)
${\mathbf{F}}_{\rm EW}(\boldsymbol{\xi})$ $=$ $E\{{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\Theta})\}=\int_{\boldsymbol{\Theta}}{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\theta})Q(d\boldsymbol{\theta})$ (S.3)
@@
Following Fedorov and Leonov (2014), we further denote the corresponding collections of Fisher information matrices:
${\cal F}_{\rm SEW}({\cal X})$ $=$ $\{{\mathbf{F}}_{\rm SEW}(\boldsymbol{\xi})\mid\boldsymbol{\xi}\in\boldsymbol{\Xi}({\cal X})\}\ ,$
${\cal F}_{\rm EW}({\cal X})$ $=$ $\{{\mathbf{F}}_{\rm EW}(\boldsymbol{\xi})\mid\boldsymbol{\xi}\in\boldsymbol{\Xi}({\cal X})\}\ .$

### v2

@@ 4 Applications to Real Experiments
In this section, we use the paper feeder experiment (see Section 1 and Example 2) and a minimizing surface defects experiment (see also Example 1), both under MLMs, to illustrate the advantages gained by adopting our algorithms and the EW D-optimal designs. For examples under GLMs, please see Section S6 of the Supplementary Material.
@@ 4.1 Paper feeder experiment
In this section, we consider the motivating example mentioned in Section 1 and Example 2, the paper feeder experiment. There are eight discrete control factors and one continuous factor, the stack force, under our consideration (see Section S4 for more details). As mentioned in Section 1, instead of using two separate GLMs to model the two types of failures, misfeed and multifeed, we consider multinomial logistic models (MLM) to model all three possible outcomes, namely misfeed, normal, and multifeed, at the same time. Using Akaike information criterion (AIC, Akaike (1973)) and Bayesian information criterion (BIC, Hastie et al. (2009)), we adopt a cumulative logit model with npo as the most appropriate model for this experiment (see Table S.1 in Section S4 of the Supplementary Material).
@@
For this experiment, we compare ten different designs (see Tables 1 and 2): (1) “Original allocation”: the original design that collected 1,785 observations roughly uniformly at 183 distinct experimental settings; (2) “EW Bu-appro”: an approximate design obtained by applying Bu et al. (2020)’s EW lift-one algorithm on the original 183 distinct settings; (3) “EW Bu-exact”: an exact design obtained by applying Bu et al. (2020)’s EW exchange algorithm on the 183 distinct settings with $n=1,785$; (4) “EW Bu-grid2.5” & (5) “EW Bu-grid2.5 exact”: approximate and exact designs by applying Bu et al. (2020)’s algorithms after discretizing the range $[0,160]$ of stack force into grid points with equal space 2.5; (6) “EW ForLion”: the proposed EW D-optimal approximate design by applying Algorithm 1 to the continuous factor, stack force, ranging in $[0,160]$; (7$\sim$(10) exact designs obtained by applying Algorithm 2 with different grid levels or pace lengths ($L=0.1,0.5,1,2.5$) and $n=1,785$. As mentioned in Example 2, for illustration purpose, all designs follow the same 18 runs modified from $OA(18,2^{1}\times 3^{7})$ for the eight discrete factors, while their design points can be different in terms of the levels of the continuous factor. Since the feasible space of a cumulative logit model is not rectangular (see Example 5.2 in Bu et al. (2020) or Example 8 in Huang et al. (2025)), we bootstrap the original dataset to collect $B=100$ (due to computational intensity) samples and the corresponding estimated model parameters, denoted by $\hat{\boldsymbol{\theta}}_{1},\hat{\boldsymbol{\theta}}_{2},\ldots,\hat{\boldsymbol{\theta}}_{100}$ All the EW designs are under the sample-based EW D-optimality with respect to the 100 parameter vectors. The constructed designs are presented in Tables LABEL:tab:D-Optimal_designs_for_the_paper_feeder_experiment and LABEL:tab:EW_Forlion_optimal_designs of the Supplementary Material.
@@
In Table 1, we list the number of design points $m$, the computational time in seconds for obtaining the designs, the objective function values, and the relative efficiencies compared with the recommended EW ForLion design, that is, $(|\hat{E}\{{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\Theta})\}|/|\hat{E}\{{\mathbf{F}}(\boldsymbol{\xi}_{\rm ForLion},$ $\boldsymbol{\Theta})\}|)^{1/p}$ with $p=32$. Compared with our EW ForLion design, the original design on 183 settings only attains $63.03\%$ relative efficiency. Having applied Bu et al. (2020)’s algorithms on the same 183 settings, the relative efficiencies of EW Bu-appro and EW Bu-exact, $88.16\%$ and $88.15\%$, are not satisfactory either. When we apply Bu et al. (2020)’s algorithms on grid points with length $2.5$, it takes much longer time to find EW Bu-grid2.5 and EW Bu-grid2.5 exact designs, which also need more settings (55 and 44). The EW ForLion design and its exact designs at different grid levels require the minimum number of experimental settings, less computational time, and higher relative efficiencies.
@@
Table 2: Summary of 100 relative efficiencies against locally D-optimal designs for paper feeder experiment
@@ 4.2 Minimizing surface defects
Wu (2008) presented a study that investigated optimal parameter settings in a polysilicon deposition process for circuit manufacturing (see also Example 1). This experiment involves responses of five categories, five continuous factors, and one discrete factor (see Table S1 in the Supplementary Material of Huang et al. (2024) or Table 6 in Lukemire et al. (2022)). Lukemire et al. (2022) employed a PSO algorithm and derived a locally D-optimal approximate design with 14 support points under a cumulative logit model with po, while Huang et al. (2024) applied their ForLion algorithm and obtained a locally D-optimal approximate design with 17 design points.
In practice, an experimenter may not know the precise values of some or all parameters but often has varying degrees of insight into their actual values. To construct a robust design against parameter misspecifications, we sample $B=1000$ parameter vectors from the prior distributions listed in Table S1 of the Supplementary Material of Huang et al. (2024) for the cumulative logit model with po as follows:
@@
with $i=1,\ldots,m$, $j=1,2,3,4$, and $J=5$, where $\pi_{ij}$ is the probability that the response associated with ${\mathbf{x}}_{i}=(x_{i1},\ldots,x_{i6})^{T}$ falls into the $j$th category.
@@
To compare EW ForLion D-optimal designs with Bu et al. (2020)’s EW D-optimal designs on grid points, we discretize each of the five continuous control factors into 2 or 4 evenly distributed grid points, which leads to $642\times 2^{5}=$ 64 or $0482\times 4^{5}=2,$ 048 possible design points, respectively. Discretizations with higher number of grid points are skipped due to computational intensity. Then Bu et al. (2020)’s lift-one algorithm is applied to the discretized design region and the corresponding EW D-optimal approximate designs are constructed (EW Bu grid2 and EW Bu grid4 in Figure 1).
@@
Design Cleaning Deposition Deposition Nitrogen Silane Settling Design Cleaning Deposition Deposition Nitrogen Silane Settling
Point Method Temp. Pressure Flow Flow Time $w_{i}$ Point Method Temp. Pressure Flow Flow Time $n_{i}$
@@
Figure 1: Relative efficiency comparison among designs for 10,000 sampled parameter sets in the minimizing surface defects experiment
@@ 5 Conclusion
Supplementary Materials
The Supplementary Material includes several sections: S1 provides analytic solutions for lift-one algorithm with MLMs; S2 discusses assumptions for design region and parametric models; S3 provides proofs of main theorems; S4 displays model selection and design comparison for paper feeder experiment; S5 discusses robustness of sample-based EW designs; S6 provides two examples for GLMs and integral-based EW D-optimality.
@@
Proofs of Main Theorems
@@
Model Selection and Design Comparison for Paper Feeder Experiment
@@
Robustness of Sample-based EW Designs
@@
In Bu et al. (2020), it was mentioned that (S1.1) has analytic solutions for $J\leq 5$, while no formula was provided accordingly. In this section, we provide explicit solutions to facilitate the programming for multinomial logistic models. Note that we only need the real-valued solutions locating in $(0,1)$ for our purposes.
@@
Lemma 2.
When $J=3$, (S1.1) can be rewritten as $A_{2}z^{2}+A_{1}z+A_{0}=0$ with $A_{2}=p(b_{0}-b_{1}+b_{2})$, $A_{1}=b_{1}(1+p)-2(b_{0}p+b_{2})$, and $A_{0}=b_{0}p-b_{1}$
@@
    If $A_{2}=A_{1}=A_{0}=0$, then (S1.1) has infinitely many solutions in $(0,1)$ and $f_{i}(z)\equiv f_{i}(0)\geq 0$. In this case, we may choose $z=0$ to maximize $f_{i}(z)$.
@@
    If $A_{2}=A_{1}=0$ but $A_{0}\neq 0$, then we must have $A_{0}>0$. In this case, (S1.1) has no solution, while $z=0$ uniquely maximizes $f_{i}(z)$ with $z\in[0,1]$.
@@
    If $A_{2}=0$ but $A_{1}\neq 0$, then (S1.1) has a unique solution $-A_{0}/A_{1}\in\mathbb{R}$.
@@
    If $A_{2}\neq 0$ and $A_{1}^{2}-4A_{2}A_{0}>0$, then (S1.1) has two real solutions
@@
    If $A_{2}\neq 0$ and $A_{1}^{2}-4A_{2}A_{0}=0$, then (S1.1) has only one real solution $-A_{1}/(2A_{2})$.
@@
    If $A_{2}\neq 0$ and $A_{1}^{2}-4A_{2}A_{0}<0$, then (S1.1) has no real solution.
Proof of Lemma 2: When $J=3$, (S1.1) can be rewritten as $A_{2}z^{2}+A_{1}z+A_{0}=0$ with $A_{2}=p(b_{0}-b_{1}+b_{2})$, $A_{1}=b_{1}(1+p)-2(b_{0}p+b_{2})$, $A_{0}=b_{0}p-b_{1}$ , and
@@
Lemma 3.
When $J=4$, equation (S1.1) is equivalent to $A_{3}z^{3}+A_{2}z^{2}+A_{1}z+A_{0}=0$ with $A_{0}=b_{0}p-b_{1}$ , $A_{1}=-3b_{0}p+b_{1}(2+p)-2b_{2}$ , $A_{2}=3b_{0}p-b_{1}(1+2p)+b_{2}(2+p)-3b_{3}$ , and $A_{3}=p(-b_{0}+b_{1}-b_{2}+b_{3})$. The cases with $A_{3}=0$ has been listed in Lemma 2. If $A_{3}\neq 0$, (S1.1) is equivalent to $z^{3}+a_{2}z^{2}+a_{1}z+a_{0}=0$ with $a_{i}=A_{i}/A_{3}$, $i=0,1,2$. We let
@@
    If $q^{3}+r^{2}>0$, then (S1.1) has only one real solution $s_{1}+s_{2}-a_{2}/3$
@@
    If $q^{3}+r^{2}=0$, then (S1.1) has two real solutions $z_{1}=2r^{1/3}-a_{2}/3$ and $z_{2}=-r^{1/3}-a_{2}/3$
@@
    If $q^{3}+r^{2}<0$, then (S1.1) has three real solutions
@@
Lemma 4.
When $J=5$, equation (S1.1) is equivalent to $A_{4}z^{4}+A_{3}z^{3}+A_{2}z^{2}+A_{1}z+A_{0}=0$ with $A_{0}=b_{0}p-b_{1}$ , $A_{1}=-4b_{0}p+b_{1}(3+p)-2b_{2}$ , $A_{2}=6b_{0}p-3b_{1}(1+p)+b_{2}(4+p)-3b_{3}$ , $A_{3}=-4b_{0}p+b_{1}(1+3p)-2b_{2}(1+p)+b_{3}(3+p)-4b_{4}$ , and $A_{4}=p(b_{0}-b_{1}+b_{2}-b_{3}+b_{4})$. The cases with $A_{4}=0$ has been listed in Lemmas 3 & 2. If $A_{4}\neq 0$, (S1.1) is equivalent to $z^{4}+a_{3}z^{3}+a_{2}z^{2}+a_{1}z+a_{0}=0$ with $a_{i}=A_{i}/A_{4}$, $i=0,1,2,3$. Then there are four solutions to (S1.1) calculated as complex numbers:
@@
Following Fedorov and Leonov (2014) and Huang et al. (2024), in this section we discuss the assumptions needed for this paper in details. Note that our assumptions have been adjusted for mixed factors. Recall that a design point or experimental setting ${\mathbf{x}}=(x_{1},\ldots,x_{d})^{T}\in{\cal X}\subseteq\mathbb{R}^{d}$. Among the $d\geq 1$ factors, the first $k$ factors are continuous, and the last $d-k$ factors are discrete. Note that $0\leq k\leq d$ and $k=0$ implies no continuous factor. When $k\geq 1$, we denote ${\mathbf{x}}_{(1)}=(x_{1},\ldots,x_{k})^{T}$ as the vector of continuous factors.
Lemma 5.
@@
The proof of Lemma 5 is straightforward. It is for sample-based EW designs. For integral-based EW designs, we need the following lemma:
Lemma 6.
@@
For any $s,t\in\{1,\ldots,p\}$, since $F_{st}(\mathbf{x},\boldsymbol{\theta})$ is continuous with respect to $\mathbf{x}_{(1)}=(x_{1},\ldots,x_{k})^{T}$, we must have $\lim_{n\rightarrow\infty}F_{st}({\mathbf{x}}_{n},\boldsymbol{\theta})=F_{st}({\mathbf{x}}_{0},\boldsymbol{\theta})$ for each $\boldsymbol{\theta}\in\boldsymbol{\Theta}$. Since (A3) is satisfied, according to the Dominated Convergence Theorem (DCT, see, for example, Theorem 5.3.3 in Resnick (2003)), we must have
$\lim_{n\rightarrow\infty}E\{F_{st}({\mathbf{x}}_{n},\boldsymbol{\Theta})\}$ $=$ $\lim_{n\rightarrow\infty}\int_{\boldsymbol{\Theta}}F_{st}({\mathbf{x}}_{n},\boldsymbol{\theta})Q(d\boldsymbol{\theta})$
@@
$=$ $E\{F_{st}({\mathbf{x}}_{0},\boldsymbol{\Theta})\}$
for each ${\mathbf{x}}_{0}\in{\cal X}$, as long as $\lim_{n\rightarrow\infty}{\mathbf{x}}_{n}={\mathbf{x}}_{0}$ That is, $E\{{\mathbf{F}}({\mathbf{x}},\boldsymbol{\Theta})\}$ is element-wise continuous with respect to ${\mathbf{x}}_{(1)}$ ∎